OpenStudy (mrnebula):

Some one help? simplify the radical expression. sqrt (42a^8)

OpenStudy (chantysquirrel1129**):

I got \[a ^{4}\sqrt{42}\]........don't know if that's what you wanted. :p

OpenStudy (green_1):

a^4sqrt(42)

OpenStudy (mrnebula):

Thats not one of the answers. plus I need some one to explain so I can do this stuff my self

OpenStudy (green_1):

@ganeshie8

OpenStudy (chantysquirrel1129**):

Well this is how: \[\sqrt{42a ^{8}}= \sqrt{42} \sqrt{a ^{8}}\] right? so \[\sqrt{4^{8}}= a ^{4}\] and put it together

OpenStudy (chantysquirrel1129**):

Sorry , I meant \[\sqrt{a ^{8}}\] not \[\sqrt{4^{8}}\]

OpenStudy (chantysquirrel1129**):

Put them together, and you get: \[a ^{4}\sqrt{42}\]

OpenStudy (chantysquirrel1129**):

Like this: \[a \sqrt{24.5}\] \[7a ^{8}\] \[7a ^{4}\] \[7\sqrt{5a ^{8}}\] and it's not that I got the answer wrong, there's just many ways to simplify it :p

OpenStudy (chantysquirrel1129**):

are those the correct answer choices or nope? :3

OpenStudy (mrnebula):

Yes

OpenStudy (mrnebula):

So is it A then? cuz idk what \[a ^{4}\sqrt{42}\] is

OpenStudy (chantysquirrel1129**):

Well this is different... ok let me think. This question has weird answer choices. My best guess would be D :o